OpenStudy (anonymous):

The measures of two complementary angles are 16z-9 and 4z + 3. Find the measures of the angles.

OpenStudy (anonymous):

By definition, two complementary angles must add up to 90. Therefore, we have

OpenStudy (anonymous):

16z-9+4z+3=90.

OpenStudy (anonymous):

do you get a decimal? i tried to solve, i got 20z-6=90 i added 6 to both sides and got 20z=96 then z =4.8?

OpenStudy (anonymous):

i'm not sure if it is supposed to be a decimal

OpenStudy (anonymous):

We combine like terms and get 20z-6=90

OpenStudy (anonymous):

Add 6 to both sides to get 20z=96 and now divide both sides by 20 gives us z=96 / 20.

OpenStudy (anonymous):

oh ok... so it is a decimaL? thanks i was confused on that

OpenStudy (anonymous):

Now using substitution, in the first equation we get (16*96/20)-9 which reduces to 67.8.This means the second angle must be 90-67.8, or 22.2, degrees.

OpenStudy (anonymous):

i like u

OpenStudy (anonymous):

We can verify by substituting 96 / 20 in the second equation: (4 * 96 / 20 ) + 3

OpenStudy (anonymous):

simplifies to 22.2.

OpenStudy (anonymous):

The Final answer is 22.2.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Hope u got it now!!! Cheers :-)!!!